OpenStudy (anonymous):

Solve for x: -3|x + 7| = -12 1)x = 5/3 , x = - 19/3 2)x = -3, x = -11 3)x = -3, x = 11 4)No solution

Vocaloid (vocaloid):

divide both sides by (-3) first, which gives us |x + 7| = 4 the absolute value sign means that there are two cases: (x+7) = 4 and (x+7) = -4 solve each of those equations for x

OpenStudy (anonymous):

What about this one? Solve for x: |x + 2| + 16 = 14 x = -32 and x = -4 x = -4 and x = 0 x = 0 and x = 28 No solution

Vocaloid (vocaloid):

after we subtract 16 from both sides we get |x+2| = -2 since absolute value is never negative, this has no solution

OpenStudy (lynfran):

the absolute values means the positive value...i.e 

OpenStudy (anonymous):

What about this one? Solve for x: -3|2x + 6| = -12 1) x = 1 and x = 5 2) x = -1 and x = -5 3) x = -9 and x = 3 4) No solution

OpenStudy (lynfran):

so x+2+16=14 x+18=14 x=14-18 x=?

OpenStudy (anonymous):

would it be no solution?

OpenStudy (lynfran):

no

OpenStudy (anonymous):

wait what problem are you talking about? I am talking about this problem Solve for x: -3|2x + 6| = -12 x = 1 and x = 5 x = -1 and x = -5 x = -9 and x = 3 No solution

OpenStudy (lynfran):

x=-4

OpenStudy (lynfran):

this one Solve for x: |x + 2| + 16 = 14